Preface

Welcome to ICNNAI’2010, the sixth International Conference on Neural Networks and Artificial Intelligence at the Brest State Technical University. ICNNAI’2010 is held in June 1-4, in Brest, Belarus.

It is hosted and organized by Ministry of Education of Republic of Belarus, Brest State Technical University and Special Interest Group of International Neural Networks Society (Belarus SIG of INNS).

Also it is in close collaboration with the Belarus State University of Informatics and Radioelectronics, United Institute of Informatics Problems of the National Academy of Sciences.

Nowadays success in Artificial Intelligence is intimately related to progress in intelligent information technologies and computer science, in particular, in the areas of neural networks, machine learning, evolutionary programming, artificial immune systems, etc., in order to deal with very complex, high-dimensional, and noisy data.

The conference aims are to present and discuss together with researchers all over the world the research results and their applications in the broad field of neural computations, introduce to some of the contemporary problems and state-of-the-art solutions at the Neural Networks and Artificial Intelligence, as well as to make new state of the art emerge by reviewing our present perspectives.

It is intended to support a worldwide exchange of knowledge’s and ideas and to faster dialogues among researchers.

The contributions to the ICNNAI’2006 cover fundamental and applied aspects in the broad field of neural computations.

After reviewing, the International Program Committee has accepted 32 submissions. Each of the accepted papers was reviewed by at least two referees. We would like to thank the referees for this painful task, which was the most important step in the selection process.

In addition to those great contributions to the regular oral sessions, the program includes two plenary talks and one mini workshop on the hottest topics of the field, all by the world top level researchers.

We also would like to thank all those who contributed to the organization of this conference.

On behalf of the Organizing Committee, we want to welcome all of the participants to ICNNAI-2010, the International Conference on Neural Networks and Artificial Intelligence 2010.

We hope this Conference will be very successful and fruitful to all of us, and will contribute to a further development in the field of Neural Networks and Artificial Intelligence. Welcome to Brest.

Chair Vladimir Golovko and Akira Imada

Co-chair Uladzimir Rubanau and Rauf Sadykhov